OpenStudy (anonymous):

How much pure antifreeze must be added to 12 gallons of 20% antifreeze to make a 80% antifreeze solution?

hero (hero):

Hint: Gallons Equation: x + 12 = y Percentage of Gallons Equation x + 12(.20) = .80y Solve the system

OpenStudy (anonymous):

x +2.4 = 80y

hero (hero):

Okay, I see what you did

hero (hero):

Can you solve the system or not?

OpenStudy (anonymous):

i dont think so, we went over how to but im kind of lost

hero (hero):

Okay, how about this, do you agree with the following mathematical statement: x = x

OpenStudy (anonymous):

yes

hero (hero):

Okay, so isolate x in both equations. Let me know what you get.

OpenStudy (anonymous):

x = 80y - 2.4 ?

hero (hero):

And what about the other equation?

hero (hero):

The two equations are related but they are still two separate equations.

OpenStudy (anonymous):

soo x = y -12?

hero (hero):

Good, now set x = x

OpenStudy (anonymous):

sorry i think this is where i get lost

hero (hero):

If x = x, that means 80y - 2.4 = y - 12 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh wait i think i got it

hero (hero):

Wait, it's supposed to be .80y - 2.4 = y - 12

hero (hero):

Don't ignore the decimals

OpenStudy (anonymous):

.2x=60 x 12

OpenStudy (anonymous):

x=36

hero (hero):

I think you did something wrong. Try again. .80y - 2.4 = y - 12

hero (hero):

Are you still working on it?

OpenStudy (anonymous):

yes, i think i see what i did

hero (hero):

@jbone, it's going to be morning by the time you figure it out.